PLease Help Me Solve this problem:A formula from anthropology says that C=100b/L. Find L if c =100 and b=9
Found 2 solutions by jim_thompson5910, fcabanski:
Answer by jim_thompson5910(35256)

C=100b/L


LC=100b


L = 100b/C


L = 100(9)/100


L = 900/100


L = 9


So the answer is L = 9


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim

Answer by fcabanski(1391)

Substitute the given values into the equation.


100=100*9/L.


Multiply both sides by L.


100L = 100*9


Divide both sides by 100.


L = (100*9)/100 = 9

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com